ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 221, 2017

Aaronson, J., M. Bromberg and H. Nakada Discrepancy skew products and affine random walks

973

Adler, R., T. Nowicki, G. Świrszcz, C. Tresser and S. Winograd Error diffusion on acute simplices: invariant tiles

445

Alvarez, S. Gibbs u-states for the foliated geodesic flow and transverse invariant measures

869

Björklund, M. and T. Hartnick Quasi-state rigidity for finitedimensional Lie algebras

25

Bou-Rabee, K. and A. Myropolska Groups with near exponential residual finiteness growth

687

Bourgain, J. and M.-C. Chang Nonlinear Roth type theorems in finite fields

853

Bowen, L. Integrable orbit equivalence rigidity for free groups  471

Bromberg, M. See J. Aaronson Celikbas, O., L. W. Christensen, L. Liang and G. Piepmeyer

Complete homology over associative rings

1

Chang, M.-C. See J. Bourgain Christensen, L. W. See O. Celikbas

Cohen, B. A generalization of the Hall–Witt identity  605

Defant, A. and A. Pérez Optimal comparison of the p-norms of Dirichlet polynomials

837

Demers, M. F. and M. Todd Equilibrium states, pressure and escape for multimodal maps with holes

367

Dick, J., A. Hinrichs, L. Markhasin and F. Pillichshammer Optimal L_{p}-discrepancy bounds for second order digital sequences

489

Dickenstein, A., M. I. Herrero and L. F. Tabera Arithmetics and combinatorics of tropical Severi varieties of univariate polynomials

741

Drusvyatskiy, D., H.-L Lee, G. Ottaviani and R. R. Thomas The euclidean distance degree of orthogonally invariant matrix varieties

291

Gao, H. Crystalline liftings and the weight part of Serre’s conjecture  117

Harpaz, Y. and M. Prasma An integral model structure and truncation theory for coherent group actions

511

Hartnick, T. See M. Björklund

Herfort, W. and W. Hojka Cotorsion and wild homology  275

Herrero, M. I. See A. Dickenstein Hinrichs, A. See J. Dick Hoffmann, D. W. and D. B. Leep Sums of squares in nonreal com-

mutative rings

803

Hojka, W. See W. Herfort

Horbez, C. The boundary of the outer space of a free product  179

Hüttemann, T. and L. Steers Finite domination and Novikov homology over strongly ℤ-graded rings

661

Idrissi, N. Swiss-Cheese operad and Drinfeld center  941

Kelmer, D. On the mean square of the remainder for the euclidean lattice point counting problem

637

Kȩpczyk, M. A ring which is a sum of two PI subrings is always a PI ring

481

Khoroshkin, S. and O. Ogievetsky Diagonal reduction algebra and the reflection equation

705

Krejčiřík, D, N. Raymond, J. Royer and P. Siegl Non-accretive Schrödinger operators and exponential decay of their eigenfunctions

779

Lazar, E. A. and R. Pemantle Coarsening in one dimension: invariant and asymptotic states

59

Lee, H.-L. See D. Drusvyatskiy Leep, D. B. See D. W. Hoffmann Liang, L. See O. Celikbas Loiskekoski, L. and G. M. Ziegler Simple polytopes without small

separators

731

Markhasin, L. See J. Dick Martínez-Cervantes, G. On weakly Radon–Nikodým compact spa-

ces

165

Myropolska, A. See K. Bou-Rabee Nakada, H. See J. Aaronson Nies, A. and K. Tent Describing finite groups by short first-order

sentences

85

Nowicki, T. See R. Adler Ogievetsky, O. See S. Khoroshkin Ottaviani, G. See D. Drusvyatskiy Pemantle, R. See E. A. Lazar Pérez, A. See A. Defant Piepmeyer, G. See O. Celikbas Pillichshammer, F. See J. Dick Prasma, M. See Y. Harpaz Raymond, N. See D. Krejčiřík Royer, J. See D. Krejčiřík

Seo, S. On the towers of torsion Bertrandias and Payan modules  563

Siegl, P. See D. Krejčiřík Soukup, D. T. Decompositions of edge-coloured infinite complete

graphs into monochromatic paths II

235

Steers, L. See T. Hüttemann Świrszcz, G. See R. Adler Tabera, L. F. See A. Dickenstein

Taylor, J. A note on skew characters of symmetric groups  435
Temkin, M. Altered local uniformization of Berkovich spaces  585

Tent, K. See A. Nies Thomas, R. R. See D. Drusvyatskiy Todd, M. See M. F. Demers Tresser, C. See R. Adler

Warnke, L. Upper tails for arithmetic progressions in random subsets  317
Weigel, T. and P. A. Zalesskii Virtually free pro-p products  425

Winograd, S. See R. Adler Zalesskii, P. A. See T. Weigel Ziegler, G. M. See L. Loiskekoski